OpenStudy (anonymous):

cos((pi/2) + x) = -sin x

OpenStudy (anonymous):

verify the identity

OpenStudy (freckles):

use the sum identity for cosine

OpenStudy (freckles):

cos(a+b)=cos(a)cos(b)-sin(a)sin(b)

OpenStudy (anonymous):

so I got cos(x+pi/2) = cosx cos(pi/2) - sin x sin(pi/2) where sin(pi/2) is then equal to 1, but what happens to the rest of the equation?

OpenStudy (freckles):

cos(pi/2)=0

OpenStudy (freckles):

so you have exactly what you need and want

OpenStudy (anonymous):

Thanks so much!